OpenStudy (anonymous):

whats the anti derivative of 1/(x^3)?? i have an answer but shows its wrong//// help?

OpenStudy (anonymous):

my answer is x^-2/2

OpenStudy (anonymous):

When doing integrals with a power on the bottom, its often helpful to rewrite it like so anti derivative of x^-3 So lets use the general rule for integration, -3 + 1 = -2 So the anti derivative would be x^(-2) / -2 or in proper form 1 / 2*x^2

OpenStudy (anonymous):

I'm sorry, I mistyped that. The answer is -2 / x^2

OpenStudy (amistre64):

x^-3 -> x^-4/4: 1/(4x^4)

OpenStudy (anonymous):

Wait I keep messing that up.... 1 / -2*x^2

OpenStudy (amistre64):

hah!!... i missed it too lol

OpenStudy (amistre64):

-1/2x^2 is right ;)

OpenStudy (anonymous):

thanks you haha ... so how did you get that>?

OpenStudy (amistre64):

its the reverse of the power rule of a derivative.

OpenStudy (anonymous):

got it .. thanks

OpenStudy (amistre64):

\[D(X^n) = n * X^{n-1} \rightarrow D(x^5) = 5x^4\]

OpenStudy (dumbcow):

you do it correctly you just forgot to include the negative when you divided by the exponent

OpenStudy (amistre64):

\[\int\limits_{} X^n dx \rightarrow \frac{X^{n+1}}{n+1}\]

OpenStudy (anonymous):

ahh i got it ,.,.,.

OpenStudy (anonymous):

thanks